# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

### v3

@@ 3 Asymptotic Theory
  • (i) The profile likelihood function $l(\theta)$ is twice differentiable with respect to $\theta\in\Theta,$ where $\Theta$ is a compact subset of $\mathbb{R}^{K}$ with $\theta_{0}$ in its interior.
  • (ii) For all $\theta\in\Theta$ and $\theta\neq\theta_{0},$ $\mathbb{P}\big{(}l(\theta)\neq l(\theta_{0})\big{)}>0.$
@@
    $\mathbb{E}\big{[}\sup_{\theta\in\Theta}|l(\theta)|\big{]}<\infty.$
  • (iv) For large $N$ and $T,$$\{\|\triangledown_{\theta}l(\theta)\|,\theta\in\Theta\}$ is uniformly integrable, and
    $\mathbb{E}\big{[}\sup_{\theta\in\Theta}\|\triangledown_{\theta\theta^{\prime}}l(\theta)\|\big{]}<\infty$
    where $\triangledown_{\theta}$ and $\triangledown_{\theta\theta^{\prime}}$ denote the first and second total derivatives with respect to $\theta.$ Here, $\|\triangledown_{\theta\theta^{\prime}}l(\theta)\|$ defines the Euclidean norm of $\triangledown_{\theta\theta^{\prime}}l(\theta).$
  • (v) The Hessian matrix $\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]$ is nonsingular for large $N$ and $T$.
Assumption 2 requires the differentiability likelihood on a compact parameter space of $\theta.$ Assumption 2 serves as an identification condition for $\theta$ ensuring that $\mathbb{E}[l(\theta)]$ is uniquely maximized at $\theta_{0}.$ Assumption 2 shows a dominance condition implying that $1/(NT)\sum_{i=1}^{N}\sum_{t=1}^{T}l_{i,t}(\theta,\phi(\theta))$ converges uniformly in probability to $\mathbb{E}[l(\theta)]$ over $\theta.$ Consequently, Assumptions 2-2 guarantee that $\widehat{\theta}_{L}$ is an asymptotically consistent estimator of $\theta_{0}.$ Assumption 2 establishes a dominance condition for the Hessian, so that for any consistent estimator $\widetilde{\theta}$ of $\theta_{0},$
$1/(NT)\sum_{i=1}^{N}\sum_{t=1}^{T}l_{i,t}(\widetilde{\theta},\phi(\widetilde{\theta}))\stackrel{{\scriptstyle\mathbb{P}}}{{\longrightarrow}}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})].$
Assumption 2 ensures that the Hessian for $\theta$ is invertible.
@@
Under Assumptions 1 and 2, we have
$\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})\stackrel{{\scriptstyle\mathbb{D}}}{{\longrightarrow}}\mathcal{N}(0,\Sigma)$
as $N$ and $T\to\infty,$ where $\Sigma=-\lim_{N,T\to\infty}\big{(}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1},$ and $\mathcal{N}(0,\Sigma)$ denotes the normal distribution with mean 0 and variance $\Sigma.$
We consider a generic nonlinear null hypothesis $H_{0}:R(\theta_{0})=0,$ where $R(\theta)$ is an $r\times 1$ vector-valued function independent of the data for some $r\leq K.$ Assume that $R(\theta)$ is continuously differentiable, with its Jacobian denoted as $J(\theta):=\triangledown_{\theta^{\prime}}R(\theta),$ and that $\operatorname*{rank}[J(\theta)]=r.$ Let $\widehat{\theta}_{R}:=\arg\max_{\theta}L(\theta)$ subject to $R(\theta)=0.$ Define the test statistics:
$\widehat{\xi}_{LR}:=-2NT\big{\{}L(\widehat{\theta}_{R})-L(\widehat{\theta}_{L})\big{\}},$
$\widehat{\xi}_{LM}:=-NT\big{\{}\triangledown_{\theta^{\prime}}L(\widehat{\theta}_{R})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{R})]^{-1}\triangledown_{\theta}L(\widehat{\theta}_{R})\big{\}},$
$\widehat{\xi}_{Wald}:=-NT\big{\{}R(\widehat{\theta}_{L})^{\prime}\big{(}J(\widehat{\theta}_{L})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}J(\widehat{\theta}_{L})^{\prime}\big{)}^{-1}R(\widehat{\theta}_{L})\big{\}}.$
The following corollary states that the hypothesis testing based on the corrected likelihood aligns with standard results in the context of MLE.
@@
Suppose Assumptions 1 and 2. Under the null hypothesis $H_{0}:R(\theta_{0})=0,$ the statistics $\widehat{\xi}_{LR},$ $\widehat{\xi}_{LM},$ and $\widehat{\xi}_{Wald}$ are asymptotically equivalent, following a common limiting $\chi^{2}(r)$ distribution with degrees of freedom $r$.
4 Simulations
In the simulation study, we consider data generating process (DGP) from a dynamic logit model as follows:
@@ 4 Simulations
where the disturbance term $\varepsilon_{it}$ follows a standard logistic distribution, independent across both $i$ and $t,$ and also independent of the regressor $X_{it}.$ According to the identification restriction, we generate true values for the incidental parameters $\{\alpha_{1,i}^{0},\gamma_{1,t}^{0}\}$ from $\mathcal{N}(0,0.01)$, after demeaning. $\{\alpha_{2,i}^{0},\gamma_{2,t}^{0}\}$ are also generated from $\mathcal{N}(0,0.01)$ where we only demean $\gamma_{2,t}^{0}$. Given these parameters, $X_{it}$ is independently generated from $\mathcal{N}\big{(}\sum_{k=1}^{2}(\alpha_{k,i}^{0}+\gamma_{k,t}^{0}),0.25\big{)}$ across both $i$ and $t.$ The true structural parameter is set as $\theta_{0}=(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}.$ The initial $Y_{i0}=\mathds{1}\{(\beta_{0}+\alpha_{1,i}^{0})X_{i0}+\alpha_{2,i}^{0}+\varepsilon_{i0}>0\}$ with $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$ and $\varepsilon_{i0}$ following a standard logistic distribution. Model (4.1) also encompasses the static logit model as a specific case where $\rho\equiv 0.$ We simulate this static case with the same DGP as the dynamic model, except for $\rho_{0}=0$ ($\theta_{0}=\beta_{0}$). We demonstrate the finite sample performance of the MLE of $\theta_{0}$ by maximizing our bias-corrected likelihood $L(\theta),$ and compare it with the uncorrected (original) MLE obtained from $\widehat{l}(\theta)$ without bias correction. In addition, we compare hypothesis testing in terms of size and power based on both the corrected and uncorrected likelihoods. We consider the sample size $(N,T)\in\{(30,30),(60,60),(90,90)\},$ with 1000 replications for each DGP.
Table 1 reports the estimation bias and RMSE of our bias-corrected and the uncorrected MLEs for both the static ($\theta_{0}=\beta_{0}$) and dynamic ($\theta_{0}=(\rho_{0},\beta_{0})^{\prime}$) logit models. In the static model, $\widehat{\beta}$ represents the uncorrected MLE from $\widehat{l}(\theta)$, and $\widehat{\beta}_{L}$ is our bias-corrected MLE. For the dynamic model, $(\widehat{\beta},\widehat{\rho})$ denotes the uncorrected MLEs, while $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ represents our bias-corrected estimates, with the truncation parameter $\tau$ set to $\tau=1$ or $2$. In the static model, ${Y_{it}}$ are generated independently across time $t$ where we set $\tau=0$ for the estimation, while the data in the dynamic model exhibit serial dependence.
Table 1: Comparison between Bias-corrected and Uncorrected MLEs
@@
    Notes: The true structure parameters are $\theta_{0}=(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $\theta_{0}=\beta_{0}=0.5$ for the static model. In the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ represent the uncorrected MLE and our bias-corrected MLE, respectively. In the dynamic model, $(\widehat{\beta},\widehat{\rho})$ is the uncorrected MLE and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ represents our bias-corrected MLE with the truncation parameter $\tau$ set to $\tau=1$ or $2$. The number of replications is $1000$.
The results in Table 1 reveal that the uncorrected MLEs are severely biased in both models, while our bias-corrected estimators substantially reduce the bias. For instance, in the static model with $(N,T)=(30,30)$, the bias for $\widehat{\beta}_{L}$ is only $0.05$, which is less than $1/3$ of the significant bias of $0.158$ for $\widehat{\beta}$. As the sample size increases to $(N,T)=(90,90)$, $\widehat{\beta}_{L}$ is nearly unbiased, while the bias for $\widehat{\beta}$ remaining five times larger than that for $\widehat{\beta}_{L}$. Furthermore, the RMSEs form our corrected MLEs are consistently lower than those of the uncorrected MLEs, especially with smaller sample sizes, underscoring the advantage of our bias correction method. Additionally, in the dynamic model, the choice of $\tau$ has little impact on the performance of our bias-corrected estimators.
We also conduct LR tests using both bias-corrected and uncorrected likelihoods. Consider the null hypothesis $H_{0}:\theta_{0}=\theta_{0}+\delta$ with $\delta\in\{0,\pm 0.2,\pm 0.15,\pm 0.1\}.$ We report the empirical size ($\delta=0$) and power ($\delta\neq 0$) at a $5\%$ significance level. Table 2 presents the empirical size based on corrected and uncorrected likelihoods, where the corrected likelihoods use truncation $\tau=0$ for the static model and $\tau=1$ for the dynamic model. The results show that the LR tests using uncorrected likelihoods suffer from severe size distortion, while the empirical size from our bias-corrected likelihoods maintains the significance level as the sample size increases. Additionally, our bias-corrected tests significantly reduce Type-I error even for small sample size.
Table 2: Empirical Size of LR Tests Based on Bias-corrected and Uncorrected Likelihoods
@@
    Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table 3 presents the empirical power of LR tests based on our bias-corrected likelihoods. The empirical power increases as both the value of $\delta$ and the sample size grow. For $(N,T)=(90,90)$, when $\delta=\pm 0.2$, the rejection rates reach nearly 100% for both the static and dynamic models. We also perform LM and Wald tests, reporting empirical size and power under the same DGPs in the appendix. The results align with those of the LR tests. Additionally, we perform simulations using probit models with results provided in the appendix.
Table 3: Empirical Power of LR Tests Based on Bias-corrected Likelihoods

### v4

@@ 3 Bias Correction and Asymptotic Theory
  • (i) The likelihood $l_{i,t}(\theta,\phi)$ is four-time continuously differentiable w.r.t. to both $\theta\in\Theta$ and $\phi\in\Phi$ almost surely, such that
    $\sup_{\theta\in\Theta}\sup_{\phi\in\Phi}\left|\partial_{\nu_{1}\cdots\nu_{S}}l_{i,t}(\theta,\phi)\right|<g(w_{it}),$
    where $g(w_{it})$ is defined in Assumption 1, $\nu_{s}$ represents any element of $(\theta^{\prime},\phi^{\prime})^{\prime}$, and $S\in\{1,2,3,4\}$.
  • (ii) For every $\theta\in\Theta$ and $\theta\neq\theta_{0}$, $\mathbb{P}[l(\theta)\neq l(\theta_{0})]>0$.
@@
    $\mathbb{E}\sup_{\theta\in\Theta}\left|l(\theta)\right|<\infty.$
  • (iv) The sequence of functions $\{\left\|\nabla_{\theta}l(\theta)\right\|:\theta\in\Theta\}$ is uniformly integrable, and
    $\mathbb{E}\sup_{\theta\in\Theta}\left\|\nabla_{\theta\theta^{\prime}}l(\theta)\right\|<\infty$
[…]
  • (v) The profiled Hessian $\mathbb{E}\nabla_{\theta\theta^{\prime}}l(\theta)$ is strictly negative definite for large $N$ and $T$.
Here Assumption 2 is similar to Assumption 1, except that we impose that $l_{i,t}(\theta,\phi)$ is four-time differentiable, and require that the derivatives of $l_{i,t}(\theta,\phi)$ w.r.t. to both $\theta$ and $\phi$ to be uniformly dominated by $g(w_{it})$. This assumption ensures that the corrected likelihood $L(\theta)$, which depends on second derivatives of $l_{i,t}(\theta,\phi)$, is twice differentiable. This is required by standard techniques used to show, e.g., the asymptotic normality of $\widehat{\theta}_{L}$. See., e.g., Hayashi (2000). Similar to Assumption 1, Assumption 2 is an identification condition for $\theta_{0}$. It guarantees that $\operatorname*{plim}_{N,T\rightarrow\infty}l(\theta)$ is uniquely maximized at $\theta_{0}$. Assumption 2 is a dominance condition implying that $l(\theta)-\mathbb{E}l(\theta)$ converges to 00 uniformly as $N,T\rightarrow\infty$. Together, Assumptions 2 to 2 are used to establish the consistency of $\widehat{\theta}_{L}$. In addition to these, Assumption 2 establishes a dominance condition for the profiled Hessian, so that $\triangledown_{\theta\theta^{\prime}}l(\overline{\theta})\overset{\mathbb{P}}{\longrightarrow}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]$ for any consistent estimator $\overline{\theta}$ of $\theta_{0}$. Assumption 2 requires that the expected profiled Hessian is strictly negative definite for sufficiently large $N$ and $T$. Assumptions 2 and 2 are additionally imposed to establish the asymptotic distributions for the quantities of interests.
The corollary below states the asymptotic properties of $\widehat{\theta}_{L}$.
@@
Under Assumptions 1 and 2, $\widehat{\theta}_{L}\overset{\mathbb{P}}{\rightarrow}\theta_{0}$ and
$\sqrt{NT}[-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]^{1/2}(\widehat{\theta}_{L}-\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}(0,\mathbb{I}_{K})$
as $N,T\rightarrow\infty$.
Here, $[-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]^{1/2}$ is the square root of $-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ which can be obtained by square-rooting the eigenvalues in the eigendecomposition of $-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$. By Assumption 2, $\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ is strictly negative definite so that $-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ only has strictly positive eigenvalues. In addition, since we focus on maximum likelihood models, we use the information matrix equality to simplify the presentation of Corollary 1. In case the information matrix equality does not hold [^6] The information matrix equality generally does not hold for, e.g., non-likelihood M-estimation models, misspecified models, and quasi-likelihood models., $-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ shall be replaced by its sandwich form and $[-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]^{1/2}$ shall be constructed thereof. See, e.g., White (1982) for detailed discussions.
For inference procedures, we consider a generic null hypothesis $H_{0}:R(\theta_{0})=0$, where $R(\theta)$ is a known $r\times 1$ ($r\leq K$) vector-valued non-random function with Jacobian $J(\theta):=\partial_{\theta^{\prime}}R(\theta)$ satisfying $\operatorname*{rank}[J(\theta)]=r$. The corollary below gives the asymptotic distribution, under $H_{0}$, of $\widehat{\xi}_{LR}$ (the LR statistic), $\widehat{\xi}_{LM}$ (the LM statistic), and $\widehat{\xi}_{Wald}$ (the Wald statistic), defined respectively as
$\widehat{\xi}_{LR}$ $:=-2NT\{L(\widehat{\theta}_{R})-L(\widehat{\theta}_{L})\},$
$\widehat{\xi}_{LM}$ $:=-NT\{\triangledown_{\theta^{\prime}}L(\widehat{\theta}_{R})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{R})]^{-1}\triangledown_{\theta}L(\widehat{\theta}_{R})\},$
$\widehat{\xi}_{Wald}$ $:=-NT\{R^{\prime}(\widehat{\theta}_{L})[J(\widehat{\theta}_{L})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}J^{\prime}(\widehat{\theta}_{L})]^{-1}R(\widehat{\theta}_{L})\},$
where $\widehat{\theta}_{R}:=\arg\max_{\theta}L(\theta)$ subject to $R(\theta)=0$.
@@
Under Assumptions 1 to 2 and $H_{0}$,
$\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}\overset{\mathbb{D}}{\longrightarrow}\chi^{2}(r)$
as $N,T\rightarrow\infty$, where $\chi^{2}(r)$ is the $\chi^{2}$-distribution with degrees of freedom $r$.
Similarly, the test statistics $\widehat{\xi}_{LM}$ and $\widehat{\xi}_{Wald}$ are constructed under the information matrix equality. Adjustments are generally required if it does not hold. For instance, White (1982) shows that $-[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}$ involved in $\widehat{\xi}_{Wald}$ shall be replaced by its sandwich form for quasi-likelihood models.
Given Theorem 2, Corollaries 1 and 2 are established using standard techniques. Particularly, Theorem 2 guarantees $L(\theta)\rightarrow_{p}\mathbb{E}l(\theta)$, so that the consistency of $\widehat{\theta}_{L}$ follows from Newey and McFadden (1994). Theorem 2 also implies that $\nabla_{\theta}L(\theta)-\nabla_{\theta}l(\theta)$ is $o_{\mathbb{P}}(1/\sqrt{NT})$, from which the asymptotic normality is established using techniques in, e.g., Hayashi (2000). The proof of Corollary 2 also closely resembles the standard techniques in, e.g., Hayashi (2000), where it is shown that all $\widehat{\xi}_{LR}$, $\widehat{\xi}_{LM}$, and $\widehat{\xi}_{Wald}$ are equivalent, converging in probability to some $\widehat{\xi}$ which is asymptotically $\chi^{2}(r)$.
4 Simulation
We conduct the simulation study in this section. We first illustrate that our procedure is capable of reducing the bias of the MLE as well as delivering likelihood-based tests with an empirical size very close to the nominal level. Followed by this, we then consider a different design in which the MLEs and the tests suffer from severe biases and size-distortions. We demonstrate that our approach leads to dramatic improvements on the estimators and substantial reductions of the Type-I error risks of the tests, even in a severely biased setting. Finally in this section, we compare our method with the Jackknife [^7] Since there has currently been no Jackknife procedures for DN-HP models. We compared our method to the Jackknife of Fernández-Val and Weidner (2016), who consider models with fixed effects only in the intercept. and show that our method out-performs.
In Design 1, we consider a dynamic binary response model as:
@@ 4 Simulation
where the true value is $\theta_{0}=(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ and the disturbance $\varepsilon_{it}$ is standard logistic, independent and identically distributed (i.i.d.) over $(i,t)$ and independent from $X_{it}$. The true incidental parameter $\{\alpha_{1,i}^{0},\gamma_{1,t}^{0}\}$ are generated from $\mathcal{N}(0,0.01)$, and demeaned after they are generated. $\{\alpha_{2,i}^{0},\gamma_{2,t}^{0}\}$ are also generated from $\mathcal{N}(0,0.01)$ but only $\gamma_{2,t}^{0}$ is demeaned. Conditional on the generated $\alpha_{k,i}^{0}$ and $\gamma_{k,t}^{0}$ ($k=1,2$), the regressor $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$ i.i.d. over $(i,t)$ and the initial $Y_{i0}=\mathds{1}\{(\beta_{0}+\alpha_{1,i}^{0})X_{i0}+\alpha_{2,i}^{0}+\varepsilon_{i0}>0\}$ with $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$ i.i.d. over $(i,t)$. We also simulate a static model, where we set $\rho\equiv 0$ in (4.1), using data generated in the same way but setting $\rho_{0}=0$ (hence $\theta_{0}=\beta_{0}$).
Table 1 reports the estimation bias and RMSE, under Design 1, of $\widehat{\theta}_{L}$ and $\widehat{\theta}$ for both the static ($\theta_{0}=\beta_{0}$) and dynamic ($\theta_{0}=(\rho_{0},\beta_{0})^{\prime}$) logit models. For the static model, $\widehat{\beta}$ denotes the uncorrected MLE obtained by maximing $\widehat{l}(\theta)$, and $\widehat{\beta}_{L}$ is the corrected estimator, obtained by maximing $L(\theta)$. For the dynamic model, $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimator for $(\rho_{0},\beta_{0})^{\prime}$, while $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ represents the bias-corrected estimator, with the truncation parameter $\tau\in(1,2)$. For the static model (with $\rho_{0}\equiv 0$), we set $\tau=0$ since the data are independent over $i$, conditional on $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}$. The results reveal that the uncorrected estimators are severely biased in both models, while our bias-corrected estimators substantially reduce the bias. For instance, in the static model with $(N,T)=(30,30)$, the bias for $\widehat{\beta}_{L}$ is only $0.05$, which is less than $1/3$ of the significant bias of $0.158$ for $\widehat{\beta}$. As the sample size increases to $(N,T)=(90,90)$, $\widehat{\beta}_{L}$ is nearly unbiased, while the bias of $\widehat{\beta}$ remaining five times larger than that of $\widehat{\beta}_{L}$. Furthermore, the RMSEs from our corrected estimators are consistently lower than those of the uncorrected estimators, especially with smaller sample sizes. This indicates that our procedure is capable of reducing the IPP bias without increasing the RMSE. For large $(N,T)$, the choice of $\tau$ has negligible impact on the performance of our bias-corrected estimators.
Table 1: Comparisons of Uncorrected and Corrected Logit Estimates, Design 1
@@
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. For the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ are, respectively, the uncorrected and corrected estimates of $\beta_{0}=0.5$. For the dynamic model $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimate of $(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ from and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ is the corrected estimate with truncation parameter $\tau\in(1,2)$. The number of replications is $1000$.
Next, under Design 1, we conduct LR tests using both the corrected and the uncorrected likelihoods for the null hypothesis $H_{0}:\theta_{0}=(0.5,0.5)^{\prime}+\delta$ with $\delta\in\{0,\pm 0.2,\pm 0.15,\pm 0.1\}$. We report the empirical size ($\delta=0$) and the empirical power ($\delta\neq 0$) at the $5\%$ significance level. For the corrected likelihood, we set the truncation parameter $\tau=0$ for the static model and $\tau=1$ for the dynamic model. Table 2 presents the empirical size. The results show that the LR test using the uncorrected likelihood suffers from a severe size distortion. On the opposite, the empirical size from our bias-corrected likelihood maintains the significance level as the sample size increases.
Table 2: Empirical Size of LR Tests Based on Corrected and Uncorrected Likelihoods for Logit Model, Design 1
@@
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level of $5\%$. The number of replications is $1000$.
Table 3 presents the empirical power of LR tests based on our corrected likelihood under Design 1. The empirical power increases as both the value of $\delta$ and the sample size grow. For $(N,T)=(90,90)$, when $\delta=\pm 0.2$, the rejection rates reach nearly $100\%$ for both the static and dynamic models. We also perform LM and Wald test, reporting the empirical size and power under the same setting in Appendix B.1. The results align with those of the LR test. We additionally perform the simulations for a probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$ i.i.d. over $(i,t)$ and independent from $X_{it}$. The results are provided in Appendix B.2.
Table 3: Empirical Power of LR Test for Logit Model using $L(\theta)$, Design 1